# The Fellowship of the LLMs: Multi-Agent Workflows for Synthetic Preference Optimization Dataset Generation

### v4

@@ Abstract
This paper presents a novel methodology for generating synthetic Preference Optimization (PO) datasets using multi-agent workflows. We evaluate the effectiveness and potential of these workflows in automating and enhancing the dataset generation process. PO dataset generation requires two modules: (1) response evaluation, and (2) response generation. In the response evaluation module, the responses from Large Language Models (LLMs) are evaluated and ranked - a task typically carried out by human annotators that we automate using LLMs. We assess the response evaluation module in a 2 step process. In step 1, we assess LLMs as evaluators using three distinct prompting strategies. In step 2, we apply the winning prompting strategy to compare the performance of LLM-as-a-Judge, LLMs-as-a-Jury, and LLM Debate. Our evaluation shows that GPT-4o-as-a-Judge is more consistent across all datasets. For the response generation module, we use the identified LLM evaluator configuration and compare different configurations of the LLM Feedback Loop. We use the win rate to determine the best multi-agent configuration for generation. Experimenting with various configurations, we find that the LLM Feedback Loop, with Llama as the generator and Gemma as the reviewer, achieves a notable 71.8% and 73.8% win rate over single-agent Llama and Gemma, respectively. After identifying the best configurations for both modules, we generate our PO datasets using the above pipeline.
The Fellowship of the LLMs: Multi-Agent Workflows for Synthetic Preference Optimization Dataset Generation
Samee Arif1and Sualeha Farid2and Abdul Hameed Azeemi1  and Awais Athar3  and Agha Ali Raza1 1Lahore University of Management Sciences, 2University of Michigan - Ann Arbor 3EMBL European Bioinformatics Institute {samee.arif, abdul.azeemi, agha.ali.raza}@lums.edu.pk sualeha@umich.edu, awais@ebi.ac.uk
@@ 1 Introduction
In the process of constructing the dataset of human preferences, the evaluation and ranking of the outputs generated by LLMs are typically done by human annotators, who assess these outputs based on various criteria such as instruction following, helpfulness, relevance, accuracy, depth, and creativity. The PO dataset generation process is divided into two modules: response evaluation and response generation. The response evaluation module involves assessing and ranking responses generated by LLMs, while the response generation module focuses on creating responses that align with the identified preferences. This manual process, while effective, is labor-intensive, time-consuming, inconsistent, and subject to human biases. In this work, we thus ask the question, Can we use LLM agents to automate and improve response evaluation and generation for constructing preference optimization (PO) datasets?.
For the response evaluation step, we leverage LLMs as evaluators and compare several configurations including LLM-as-a-Judge, LLMs-as-a-Jury, and LLM Debate to pick the best evaluation strategy. The selected response evaluation module is used to evaluate and identify the optimal response generation module. Previously, single-agents have been used to generate the responses for PO datasets; however, we use a multi-agent framework for response generation, which allows us to generate more refined, higher-quality responses. The multi-agent approach uses the collaboration between multiple LLMs, where one agent can provide suggestions for improvements, and the other can revise the response based on the feedback. This iterative process leads to a thorough refinement of the generated content, ensuring that the final output better aligns with human preferences and expectations.
@@
  1. 1.
    We generate synthetic PO datasets for LLM improvement by combining the best configuration for the evaluation and generation module.
  2. 2.
    We present a comprehensive evaluation of using LLMs as evaluators on the task of selecting the better response among the candidate responses. We specifically compare the performance of three distinct approaches: LLM-as-a-Judge, LLMs-as-a-Jury, and LLM Debate.
  3. 3.
    We present an evaluation of the LLM Feedback Loop workflow for the response generation module, specifically testing different configurations using Llama-3.1-8 and Gemma-2-9b models.
@@ 2.1 Preference Optimization
Preference Optimization has emerged as a pivotal technique for aligning model outputs with human preferences. Rafailov et al. (2024) introduce DPO, a method that simplifies solving the standard RLHF problem by converting it into a classification task, enabling the extraction of the optimal policy in a straightforward way. Hong et al. (2024) introduce ORPO algorithm that combines the traditional supervised fine-tuning and preference alignment stages into a single process. The dataset for DPO and ORPO require annotated preference pairs, where each pair consists of two model outputs labeled according to which one better aligns with human preferences. Ethayarajh et al. (2024) introduce KTO, a cost-effective approach to align Large Language Models (LLMs) with human feedback, improving performance without the need for preference pairs. Argilla Distilabel (Álvaro Bartolomé Del Canto et al., 2024) uses LLM to judge between the responses of two models to create synthetic PO datasets. The datasets are available on Hugging Face [^1] https://huggingface.co/argilla. To our knowledge, no one has yet explored the use of Multi-Agent workflows for the generation of PO datasets.
2.2 Agent Frameworks
@@
Recently, there has been a growing interest in using LLM multi-agent frameworks for different tasks. Zheng et al. (2023a) presents an evaluation of LLM-as-a-Judge on the MT-Bench (Zheng et al., 2023b) and Chatbot Arena (Li et al., 2024). Their results reveal that strong LLM judges like GPT-4 can match both controlled and crowd-sourced human preferences well, achieving over 80% agreement, the same level of agreement between humans. Additionally, they evaluate several variants of Llama and Vicuna on the dataset. They study the limitations of LLM-as-a-judge, including position, verbosity, and self-enhancement biases, as well as limited reasoning ability. Verga et al. (2024) explore the use of LLMs-as-a-Jury. Their approach, a Panel of LLM evaluators (PoLL), composed of a larger number of smaller models outperforms a single large judge. They also show that the PoLL approach exhibits less intra-model bias as compared to LLM-as-a-Judge. They use Command-R, GPT, Claude-3, and Mistral families for their study. Additionally, they compare two prompting strategies: (1) reference-based scoring where they provide the LLM with a reference answer, and (2) candidate answer and pair-wise scoring where they ask the LLM to pick the better response from the candidate responses. PoLL outperforms single-agents on KILT (Petroni et al., 2021) and Chatbot Arena.
@@
Liang et al. (2024) introduce Multi-Agent Debate (MAD) to encourage divergent thinking in LLMs. They mitigate the Degeneration-of-Thought (DoT) problem, which is that once the LLM has established confidence in its solutions, it is unable to generate novel thoughts. In their approach, the affirmative LLM and the negative LLM debate on the answer while the LLM judge evaluates both arguments after each round of debate. They evaluate the approach on the Commonsense Machine Translation Dataset (Chinese to English) (He et al., 2020) and their Counter-Intuitive Arithmetic Reasoning (CIAR) dataset. MAD was able to achieve a 37% accuracy on the CIAR dataset using GPT-3.5-Turbo which outperforms Chain-of-Thought, Self-Consistency, and Self-Reflection prompting. They also show that using the MAD approach decreases bias and increases response diversity. Du et al. (2023) evaluates a different variant of multi-agent debate where multiple models generate their own responses, and each model receives the opinions of the other models, then updates its response if necessary. This is done for multiple rounds. Du et al. (2023) evaluates the approach on the following tasks: Biography generation, MMLU, Chess move validity and optimality, Arithmetic, and Grade school math,. Their approach using ChatGPT and Bard outperforms single-agent on all the tasks. To evaluate LLM responses Chan et al. (2023) presents another variant of multi-agent debate. Their architecture involves assigning agents different roles such as General Public, Critic, Psychologist, News Author, and Scientist. They used ChatGPT and GPT-4 for their evaluation on FairEval (Wang et al., 2023a) dataset and achieved a Cohen’s Kappa score of 0.40 using LLM Debate, 0.03 more than the single-agent.
@@ 3.1 Experimental Setup
In this study, we perform experiments on the three categories of models given in Table 1. For the evaluation module, we evaluate single-agents and multi-agent frameworks on four datasets, Alpaca Eval (Li et al., 2023), FairEval (Wang et al., 2023a), PandaLM-Eval (Wang et al., 2024, 2023b) and MT-Bench (Zheng et al., 2023b). For the generation module, we compare the multi-agent frameworks using win rate - the ratio of times a generation framework is selected as the best by an LLM evaluator when comparing outputs from all generation workflows. After the extensive evaluation of both modules, we used the picked strategies to generate synthetic PO datasets. We set the temperature to 0 in all our evaluations to ensure reproducibility.
@@ 3.2 LLM-as-Evaluator
With the aim of automating the evaluation component of PO dataset generation, we assess the performance of LLMs in the role of evaluators using the Alpaca Eval, FairEval, PandaLM-Eval, and MT-Bench datasets. Our goal is to determine whether multi-agent workflows work better than a single-agent for LLM evaluation. The system prompts for this task are modified version of the prompts used by Zheng et al. (2023a) and are given in Appendix A.
@@
For each of these prompting strategy, we systematically analyze the performance of the LLMs by calculating Cohen’s Kappa, against the human annotations. The system prompts are given in Table LABEL:tab:prompt-llm-judge in Appendix A. ko
@@
We also evaluate the LLM Debate framework following the implementation described by Chan et al. (2023). In this approach, we assign three distinct roles—Psychologist, General Public, and Critic—and the three agents debate the scores that should be assigned to candidate responses. After the debate, each agent gives its final score which is used to determine which candidate response they vote for. These votes are then used to pick the best response. This strategy is evaluated using the FairEval, PandaLM-Eval, and MT-Bench benchmarks. Figure 1 illustrates the debate workflow employed in our study. The system prompt, the user message structure and the prompts for the roles used are given in Table LABEL:tab:prompt-llm-debate and Table LABEL:tab:prompt-roles in Appendix A.
@@ 3.3 LLM-as-Generator
To evaluate the LLM Feedback Loop workflow for the generation module, we test different configurations using Llama-3.1-8b (Meta, 2024) and Gemma-2-9b (Google, 2024) models. In this framework, a generator LLM produces a response, which is then evaluated by a feedback LLM that provides improvement suggestions as shown in Figure 2. The generator revises the response based on these suggestions, and the process repeats for multiple iterations. The system prompt for the generator and reviewer is given in Table LABEL:tab:generator_prompt and LABEL:tab:reviewer_prompt in Appendix A. We calculate the win rate against single-agent GPT-4o (OpenAI, 2024), Llama-3.1-8b and Gemma-2-9b baseline outputs on a subset of 500 prompts from the Argilla Capybara DPO dataset [^2] https://huggingface.co/datasets/argilla/distilabel-capybara-dpo-7k-binarized to identify the best configuration. We test the following configuration:
@@
    Same Model as Both Agents: Gemma-2-9b or Llama-3.1-8b as both the feedback and generation agent.
@@
    Different Models for Each Agent: Gemma-2-9b as the feedback agent and Llama-3.1-8b as the generation agent, or vice versa.
@@
    Both Models for Feedback, One for Generation: Gemma-2-9b or Llama-3.1-8b as the generation agent, with both models as feedback agents.
@@ 3.4 Preference Optimization Dataset
We use the best configurations of the generation and evaluation modules to generate the DPO and KTO datasets. The generation module produces $N$ responses (where $N$ is the number of feedback iterations), which are then passed to the evaluation module. The evaluation module sorts these responses into the accepted and rejected fields in the DPO and KTO datasets. In this study, we use the prompts from the Argilla Capybara DPO dataset. The prompt templates used for LLM improvement dataset generation are given in Table LABEL:tab:prompt-llm-judge, LABEL:tab:generator_prompt and LABEL:tab:reviewer_prompt in Appendix A. The evaluation code, all the evaluation outputs and the generated datasets are publicly available on GitHub [^3] https://github.com/ulrs0/MA-PO.
@@ 4.1 LLM-as-Evaluator
Figure 3 shows that GPT-4o selects GPT-3.5-Turbo as the better agent 50 times and Vicuna-13b 30 times. This indicates a potential bias in favor of GPT responses when GPT-4o is the evaluator. Additionally, we can observe in the figure that Llama models also display a similar bias towards GPT responses, whereas Gemma models do not exhibit this bias, suggesting that Gemma is more impartial in its evaluations.
Figure 3: Number of times GPT-3.5-Turbo and Vicuna-13b are picked by each LLM Judge.
@@
The jury approach, by incorporating diverse models, mitigates the biases that occur in LLM-as-a-Judge approach (as shown in Figure 3) when bench-marking on the Fair-Eval dataset. However while the jury approach can offer robustness through diversity, in evaluation task, it does not universally outperform single judges. The decision to employ a jury versus a judge should consider whether the candidate responses being evaluated include output from the judge itself, which can introduce bias in the results. Additionally, scalability should be taken into account, as the jury approach might require more computational resources. Another critical consideration is the variability in performance across different datasets, which poses a challenge for generalization.
@@
Based on the comparative evaluation scores across the three datasets and the advantages and disadvantages associated with each multi-agent framework, we have chosen to use the LLM-as-a-Judge approach with GPT-4o as our primary evaluator for generating the PO dataset. This decision is driven by multiple factors:

### v5

# The Fellowship of the LLMs: Multi-Model Workflows for Synthetic Preference Optimization Dataset Generation

@@ Abstract
This paper presents a novel methodology for generating synthetic Preference Optimization (PO) datasets using multi-model workflows. We evaluate the effectiveness and potential of these workflows in automating and enhancing the dataset generation process. PO dataset generation requires two modules: (1) response evaluation, and (2) response generation. In the response evaluation module, the responses from Large Language Models (LLMs) are evaluated and ranked - a task typically carried out by human annotators that we automate using LLMs. We assess the response evaluation module in a 2 step process. In step 1, we assess LLMs as evaluators using three distinct prompting strategies. In step 2, we apply the winning prompting strategy to compare the performance of LLM-as-a-Judge, LLMs-as-a-Jury, and LLM Debate. Our evaluation shows that GPT-4o-as-a-Judge is more consistent across all datasets. For the response generation module, we use the identified LLM evaluator configuration and compare different configurations of the LLM Feedback Loop. We use the win rate to determine the best multi-model configuration for generation. Experimenting with various configurations, we find that the LLM Feedback Loop, with Llama as the generator and Gemma as the reviewer, achieves a notable 71.8% and 73.8% win rate over single-model Llama and Gemma, respectively. After identifying the best configurations for both modules, we generate our PO datasets using the above pipeline.
The Fellowship of the LLMs: Multi-Model Workflows for Synthetic Preference Optimization Dataset Generation
Samee Arif1*, Sualeha Farid2*, Abdul Hameed Azeemi1 Awais Athar3, Agha Ali Raza1 1Lahore University of Management Sciences, 2University of Michigan - Ann Arbor 3EMBL European Bioinformatics Institute {samee.arif, abdul.azeemi, agha.ali.raza}@lums.edu.pk sualeha@umich.edu, awais@ebi.ac.uk
11footnotetext: These authors contributed equally to this work.
@@ 1 Introduction
In the process of constructing the dataset of human preferences, the evaluation and ranking of the outputs generated by LLMs are typically done by human annotators, who assess these outputs based on various criteria such as instruction following, helpfulness, relevance, accuracy, depth, and creativity. The PO dataset generation process is divided into two modules: response evaluation and response generation. The response evaluation module involves assessing and ranking responses generated by LLMs, while the response generation module focuses on creating responses that align with the identified preferences. This manual process, while effective, is labor-intensive, time-consuming, inconsistent, and subject to human biases. In this work, we thus ask the question, Can we use LLMs to automate and improve response evaluation and generation for constructing preference optimization (PO) datasets?.
For the response evaluation step, we leverage LLMs as evaluators and compare several configurations including LLM-as-a-Judge, LLMs-as-a-Jury, and LLM Debate to pick the best evaluation strategy. The selected response evaluation module is used to evaluate and identify the optimal response generation module. Previously, single-models have been used to generate the responses for PO datasets; however, we use a multi-model framework for response generation, which allows us to generate more refined, higher-quality responses. The multi-model approach uses the collaboration between multiple LLMs, where one model can provide suggestions for improvements, and the other can revise the response based on the feedback. This iterative process leads to a thorough refinement of the generated content, ensuring that the final output better aligns with human preferences and expectations.
@@ 2.1 Preference Optimization
Preference Optimization has emerged as a pivotal technique for aligning model outputs with human preferences. Rafailov et al. (2024) introduce DPO, a method that simplifies solving the standard RLHF problem by converting it into a classification task, enabling the extraction of the optimal policy in a straightforward way. Hong et al. (2024) introduce ORPO algorithm that combines the traditional supervised fine-tuning and preference alignment stages into a single process. The dataset for DPO and ORPO require annotated preference pairs, where each pair consists of two model outputs labeled according to which one better aligns with human preferences. Ethayarajh et al. (2024) introduce KTO, a cost-effective approach to align Large Language Models (LLMs) with human feedback, improving performance without the need for preference pairs. Argilla Distilabel (Álvaro Bartolomé Del Canto et al., 2024) uses LLM to judge between the responses of two models to create synthetic PO datasets. The datasets are available on Hugging Face [^1] https://huggingface.co/argilla. To our knowledge, no one has yet explored the use of multi-model workflows for the generation of PO datasets.
2.2 Multi-Model Frameworks
@@
Recently, there has been a growing interest in using LLM multi-model frameworks for different tasks. Zheng et al. (2023a) presents an evaluation of LLM-as-a-Judge on the MT-Bench (Zheng et al., 2023b) and Chatbot Arena (Li et al., 2024). Their results reveal that strong LLM judges like GPT-4 can match both controlled and crowd-sourced human preferences well, achieving over 80% agreement, the same level of agreement between humans. Additionally, they evaluate several variants of Llama and Vicuna on the dataset. They study the limitations of LLM-as-a-judge, including position, verbosity, and self-enhancement biases, as well as limited reasoning ability. Verga et al. (2024) explore the use of LLMs-as-a-Jury. Their approach, a Panel of LLM evaluators (PoLL), composed of a larger number of smaller models outperforms a single large judge. They also show that the PoLL approach exhibits less intra-model bias as compared to LLM-as-a-Judge. They use Command-R, GPT, Claude-3, and Mistral families for their study. Additionally, they compare two prompting strategies: (1) reference-based scoring where they provide the LLM with a reference answer, and (2) candidate answer and pair-wise scoring where they ask the LLM to pick the better response from the candidate responses. PoLL outperforms single-models on KILT (Petroni et al., 2021) and Chatbot Arena.
@@
Liang et al. (2024) introduce Multi-Agent Debate (MAD) to encourage divergent thinking in LLMs. They mitigate the Degeneration-of-Thought (DoT) problem, which is that once the LLM has established confidence in its solutions, it is unable to generate novel thoughts. In their approach, the affirmative LLM and the negative LLM debate on the answer while the LLM judge evaluates both arguments after each round of debate. They evaluate the approach on the Commonsense Machine Translation Dataset (Chinese to English) (He et al., 2020) and their Counter-Intuitive Arithmetic Reasoning (CIAR) dataset. MAD was able to achieve a 37% accuracy on the CIAR dataset using GPT-3.5-Turbo which outperforms Chain-of-Thought, Self-Consistency, and Self-Reflection prompting. They also show that using the MAD approach decreases bias and increases response diversity. Du et al. (2023) evaluates a different variant of multi-model debate where multiple models generate their own responses, and each model receives the opinions of the other models, then updates its response if necessary. This is done for multiple rounds. Du et al. (2023) evaluates the approach on the following tasks: Biography generation, MMLU, Chess move validity and optimality, Arithmetic, and Grade school math,. Their approach using ChatGPT and Bard outperforms single-model on all the tasks. To evaluate LLM responses Chan et al. (2023) presents another variant of multi-model debate. Their architecture involves assigning models different roles such as General Public, Critic, Psychologist, News Author, and Scientist. They used ChatGPT and GPT-4 for their evaluation on FairEval (Wang et al., 2023a) dataset and achieved a Cohen’s Kappa score of 0.40 using LLM Debate, 0.03 more than the single-model.
@@ 3.1 Experimental Setup
In this study, we perform experiments on the three categories of LLMs given in Table 1. For the evaluation module, we evaluate single-models and multi-model frameworks on four datasets, Alpaca Eval (Li et al., 2023), FairEval (Wang et al., 2023a), PandaLM-Eval (Wang et al., 2024, 2023b) and MT-Bench (Zheng et al., 2023b). For the generation module, we compare the multi-model frameworks using win rate - the ratio of times a generation framework is selected as the best by an LLM evaluator when comparing outputs from all generation workflows. After the extensive evaluation of both modules, we used the picked strategies to generate synthetic PO datasets. We set the temperature to 0 in all our evaluations to ensure reproducibility.
@@ 3.2 LLM-as-Evaluator
With the aim of automating the evaluation component of PO dataset generation, we assess the performance of LLMs in the role of evaluators using the Alpaca Eval, FairEval, PandaLM-Eval, and MT-Bench datasets. Our goal is to determine whether multi-model workflows work better than a single-model for LLM evaluation. The system prompts for this task are modified version of the prompts used by Zheng et al. (2023a) and are given in Appendix A.
@@
For each of these prompting strategy, we systematically analyze the performance of the LLMs by calculating Cohen’s Kappa, against the human annotations. The system prompts are given in Table LABEL:tab:prompt-llm-judge in Appendix A.
@@
We also evaluate the LLM Debate framework following the implementation described by Chan et al. (2023). In this approach, we assign three distinct roles—Psychologist, General Public, and Critic—and the three models debate the scores that should be assigned to candidate responses. After the debate, each model gives its final score which is used to determine which candidate response they vote for. These votes are then used to pick the best response. This strategy is evaluated using the FairEval, PandaLM-Eval, and MT-Bench benchmarks. Figure 1 illustrates the debate workflow employed in our study. The system prompt, the user message structure and the prompts for the roles used are given in Table LABEL:tab:prompt-llm-debate and Table LABEL:tab:prompt-roles in Appendix A.
@@ 3.3 LLM-as-Generator
To evaluate the LLM Feedback Loop workflow for the generation module, we test different configurations using Llama-3.1-8b (Meta, 2024) and Gemma-2-9b (Google, 2024) models. In this framework, a generator LLM produces a response, which is then evaluated by a feedback LLM that provides improvement suggestions as shown in Figure 2. The generator revises the response based on these suggestions, and the process repeats for multiple iterations. The system prompt for the generator and reviewer is given in Table LABEL:tab:generator_prompt and LABEL:tab:reviewer_prompt in Appendix A. We calculate the win rate against single-model GPT-4o (OpenAI, 2024), Llama-3.1-8b and Gemma-2-9b baseline outputs on a subset of 500 prompts from the Argilla Capybara DPO dataset [^2] https://huggingface.co/datasets/argilla/distilabel-capybara-dpo-7k-binarized to identify the best configuration. We test the following configuration:
@@
    Same Model: Gemma-2-9b or Llama-3.1-8b as both the feedback and generation model.
@@
    Different Models: Gemma-2-9b as the feedback model and Llama-3.1-8b as the generation model, or vice versa.
@@
    Both Models for Feedback, One for Generation: Gemma-2-9b or Llama-3.1-8b as the generation model, with both models as feedback model.
@@ 3.4 Preference Optimization Dataset
We use the best configurations of the generation and evaluation modules to generate the DPO and KTO datasets. The generation module produces $N$ responses (where $N$ is the number of feedback iterations), which are then passed to the evaluation module. The evaluation module sorts these responses into the accepted and rejected fields in the DPO and KTO datasets. In this study, we use the prompts from the Argilla Capybara DPO dataset. The prompt templates used for LLM improvement dataset generation are given in Table LABEL:tab:prompt-llm-judge, LABEL:tab:generator_prompt and LABEL:tab:reviewer_prompt in Appendix A. The evaluation code, all the evaluation outputs and the generated datasets are publicly available on GitHub [^3] https://github.com/sameearif/Fellowship-of-The-LLMs.
@@ 4.1 LLM-as-Evaluator
We calculate the Agreement between the LLM evaluator and human evaluator for Vicuna-13b and GPT-3.5-Turbo separately. In the formula below, the numerator represents the number of instances where the LLM evaluator picks model A’s response over model B, while the denominator represents the total number of instances where humans labeled model A as the better response.
[…]
The Bias Score, as given below, provides insight into potential bias in the LLM evaluator. If the difference is positive with a high magnitude, it indicates a bias toward Vicuna-13b, as the evaluator aligns more closely with human preferences for Vicuna-13b. Conversely, if the difference is negative with a high magnitude, it suggests a bias toward GPT-3.5-Turbo. A small magnitude (close to zero) implies that the LLM evaluator is relatively unbiased, showing similar levels of agreement with human preferences for both models.
[…]
The Bias Score highlights potential biases in the LLM evaluator’s alignment with human preferences for Vicuna-13b and GPT-3.5-Turbo, as shown in Table 4. Bias Score for Llama-3.1-8b (+0.51) and Llama-3.1-70b (+0.37), indicates a strong bias toward Vicuna-13b, where the evaluator more frequently favors Vicuna-13b over GPT-3.5-Turbo. Conversely, for Gemma-2-9b (-0.05) and Gemma-2-27b (+0.02) the small magnitude of Bias Score suggests that Gemma models are impartial. For GPT-4o-mini (+0.18) and GPT-4o (+0.33), the Bias Score indicates a moderate bias toward Vicuna-13b, as the evaluator shows a noticeable but less pronounced preference for Vicuna-13b’s responses compared to GPT-3.5-Turbo. Vicuna-13b is fine-tuned on the ShareGPT dataset, which contains conversations from GPT-4 and GPT-3.5. This fine-tuning likely aligns Vicuna-13b’s responses with GPT models, explaining the evaluator’s bias toward it.
Model Agreement Bias Score
Vicuna-13b GPT-3.5-Turbo
Gemma-2-9b 0.68 0.73 -0.05
Llama-3.1-8b 0.92 0.41 +0.51
Gemma-2-27b 0.80 0.78 +0.02
Llama-3.1-70b 0.88 0.51 +0.37
GPT-4o-mini 0.84 0.66 +0.18
GPT-4o 0.92 0.59 +0.33
Table 4: Agreement between the LLM evaluator and human evaluator over Vicuna-13b and GPT-3.5 separately.
@@
The jury approach, by incorporating diverse models, mitigates the biases that occur in LLM-as-a-Judge approach when bench-marking on the Fair-Eval dataset. However while the jury approach can offer robustness through diversity, in evaluation task, it does not universally outperform single judges. The decision to employ a jury versus a judge should consider whether the candidate responses being evaluated include output from the judge itself, which can introduce bias in the results. Additionally, scalability should be taken into account, as the jury approach might require more computational resources. Another critical consideration is the variability in performance across different datasets, which poses a challenge for generalization.
@@
Based on the comparative evaluation scores across the three datasets and the advantages and disadvantages associated with each multi-model framework, we have chosen to use the LLM-as-a-Judge approach with GPT-4o as our primary evaluator for generating the PO dataset. This decision is driven by multiple factors: